please help me solve this: Tony is half as old as Andrew. Cindy is 6 years younger than twice Andrew's age. If the sum of Tony, Andrew, and Cindy's ages is 118, how old is Cindy? Type in Cindy's age to the nearest hundredth of a year
Answer by stanbon(75887)

Tony is half as old as Andrew. Cindy is 6 years younger than twice Andrew's age. If the sum of Tony, Andrew, and Cindy's ages is 118, how old is Cindy? Type in Cindy's age to the nearest hundredth of a year
-------
Equations:
a = 2t
c = 2a-6 = 2(2t)-6 = 4t-6
-------
t + a + c = 118
t + 2t + 4t-6 = 118
7t - 6 = 118
7t = 124
t = 17.71 years (tony's age now)
c = 4*t-6 = 64.86 yrs (Cindy's age now)
============================
Cheers,
Stan H.
============================
-----